OpenStudy (anonymous):

∑n=1∞(cosnΠ)/2n..converges or diverges?please explain!

OpenStudy (anonymous):

∑n=1∞(cosnΠ)/2n\[\sum_{1}^{\infty} (\cos n \Pi)/2^{n}\]

OpenStudy (anonymous):

cos(nPi)=(-1)^(n) so you have an alternating series that is monotone decreasing, so it should converge

OpenStudy (anonymous):

diverges

OpenStudy (watchmath):

take rsvitale answer!

OpenStudy (anonymous):

so what happens to 2^n ??

OpenStudy (anonymous):

is it /2n or /2^n ?

OpenStudy (anonymous):

we have an alternating series, so 1/2^n is the "other part" it is always positive and monotone decreasing.

OpenStudy (anonymous):

/(2n) and /(2^n) both should converge

OpenStudy (anonymous):

what would be the sum and please explain in steps!

OpenStudy (anonymous):

ok do you know about geometric series?

OpenStudy (anonymous):

yes..is 1/2^n geometric?

OpenStudy (anonymous):

we split the numerator and denominator into two sums right?

OpenStudy (anonymous):

yes, but some of the terms are negative and some are positive so we need two sums

OpenStudy (anonymous):

and then what would be the sum to infinity for cos n pi??

OpenStudy (anonymous):

no you dont split numerator and denominator

OpenStudy (anonymous):

the terms will be -1/2+1/4-1/8+1/16. we write the negative terms (n is odd) as one sum and the positive terms (n is even) as another sum and add them

OpenStudy (anonymous):

ok..and then?

OpenStudy (anonymous):

what would be the sum for entire series?

OpenStudy (anonymous):

the negative terms: \[-\frac{1}{2}\sum_{n=0}^{\infty}\frac{1}{4^n}\]

OpenStudy (anonymous):

@rsvitale: Hey! thanks ..understood!

OpenStudy (anonymous):

Is sigma n=1 to infinity of (3^n -1) divided by 2^n divergent? just tell me if it is right or wrong

OpenStudy (anonymous):

positive terms: \[\sum_{n=1}^{\infty}\frac{1}{4^n}\] you should get -1/3 as the total sum when you add the two together

OpenStudy (anonymous):

let me look for a sec

OpenStudy (anonymous):

this?: \[\sum_{n=1}^{\infty}\frac{3^n-1}{2^n}\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

that is the same as: \[\sum_{n=1}^{\infty}\left( \frac{3}{2} \right) ^n-\sum_{n=1}^{\infty}\left( \frac{1}{2} \right)^n\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

the first diverges since it is a geometric series with term greater than or equal to 1, So the whole thing diverges